OpenStudy (anonymous):

The table shows the results of a spinner with four colored sections. Use the table to answer the question. The spinner is spun 50 more times. Which is the best prediction for the number of times the spinner will stop on blue? A. 6 B. 12 C. 24 D. 48

OpenStudy (anonymous):

here is chart

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can u help me

OpenStudy (lucaz):

lets see..

OpenStudy (anonymous):

i have pic of chart

OpenStudy (anonymous):

do u got it

OpenStudy (anonymous):

do u got it yet lucaz

OpenStudy (lucaz):

notice when you add the number of times you get 100, so you can say that the chance getting blue is 24/100 or 24%

OpenStudy (anonymous):

thx u r smart

OpenStudy (lucaz):

now, I'm not sure if this is correct, but if you multiply 24% by 50 you get \[\large{\frac{24}{100}\times 50=\frac{120}{10}=12}\] I don't know if this is really the answer

OpenStudy (lucaz):

actually it seems it is correct, because for 100 times the change is 24% so for 50 times, half of 100 you get half of 24% wich is 12%, I think it's correct